OpenStudy (anonymous):

A tsunami of wavelength 230 km and velocity 675 km/h travels across the Pacific Ocean. As it approaches Hawaii people observe an unusual decrease in sea level in the harbors corresponding to the trough of the incoming wave. Approximately how much time do they have to run to safety? So far I worked out: (1/2)lambda--> (1/2)230=115km (115km/675 km/h)= .170 hrs But how would I convert this to minutes?!?! Do I multiply .170hrs by 60s?

OpenStudy (anonymous):

Proceed by multiplying by variations of 1. For instance, to convert \(X\) in hours to minutes, perform the following multiplication.\[\left( X \text{ hours} \right) \left( \frac{60 \text{ minutes}}{1 \text{ hour}} \right)\]

OpenStudy (anonymous):

Take note how the \(\text{hours}\) units cancel as if they were variables.